OpenStudy (peperwqio):

Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse? Correct answer gets a medal.

OpenStudy (anthonyym):

You can find out how many hours it take first. \[\frac{ 55mi }{ hr }=\frac{ 310mi }{ x }\]

OpenStudy (peperwqio):

So the correct anser is...?

OpenStudy (anthonyym):

\[\frac{ 55mi }{ 1 hr }=\frac{ 310mi }{ x }\] Can you solve it now through cross multiplying?

OpenStudy (peperwqio):

5.64?

OpenStudy (anthonyym):

Yes 5.64 hours. Now convert to minutes. \[\frac{ 5.64hr }{ 1 }*\frac{ 60 minutes }{1 hr }*\frac{ 60 seconds }{ 1 minute }\]. Notice the units hour and minutes cancel.

OpenStudy (anthonyym):

Oops my bad skip the last part. It asks for minutes not seconds

OpenStudy (anthonyym):

Should be \[\frac{ 5.64hr }{ 1 }*\frac{ 60 minutes }{1 hr }\]